ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 171, 2009

Bell, J. P. Division algebras of Gelfand–Kirillov transcendence degree 2

51

Berkovich, Y. and Z. Janko On subgroups of finite p-groups  29

Blanco, A. and N. Grønbæk Amenability of algebras of approximable operators

127

Brodskiy, N., J. Dydak, J. Higes and A. Mitra Assouad–Nagata dimension via Lipschitz extensions

405

Clark, P. L. On the Hasse principle for Shimura curves  349

Colombo, F., I. Sabadini and D. C. Struppa Slice monogenic functions

385

De Bernardi, C. and L. Veselý On support points and support functionals of convex sets

15

De Filippis, V. Generalized derivations with Engel condition on multilinear polynomials

325

Dubinin, V. N. and M. Vuorinen On conformal moduli of polygonal quadrilaterals

111

Dydak, J. See N. Brodskiy Džamonja, M., J. A. Larson and W. J. Mitchell A partition

theorem for a large dense linear order

237

Fishman, L. Schmidt’s game on fractals  77

van Gaans, O. and B. Lemmens Dynamics of non-expansive maps on strictly convex Banach spaces

425

Gillet, H. and C. Soulé Erratum for: “On the number of lattice points in convex symmetric bodies and their duals”, Israel Journal of Mathematics, Vol. 74, 1991, pp. 347–357

443

Grønbæk, N. See A. Blanco Higes, J. See N. Brodskiy Janko, Z. See Y. Berkovich

Keller, T. M. Counting characters in linear group actions  367

Larson, J. A. See M. Džamonja Lemmens, B. See O. van Gaans Maříková, J. The structure on the real field generated by the stan-

dard part map on an o-minimal expansion of a real closed field 175 Matsutani, S. and E. Previato A generalized Kiepert formula for

C_{ab} curves 305

Mattarei, S. A sufficient condition for a number to be the order of a nonsingular derivation of a Lie algebra

1

Mitchell, W. J. See M. Džamonja Mitra, A. See N. Brodskiy

Nair, R. On pair correlations and Hausdorff dimension  197
Navarro, G., L. Sanus and P. H. Tiep Real characters and degrees  157
Nilsson, J. On numbers badly approximable by dyadic rationals  93

Pina, R. and K. Tenenblat On solutions of the Ricci curvature equation and the Einstein equation

61

Previato, E. See S. Matsutani Ruscheweyh, S., L. Salinas and T. Sugawa Completely mono-

tone sequences and universally prestarlike functions

285

Sabadini, I. See F. Colombo Salinas, L. See S. Ruscheweyh Sanus, L. See G. Navarro Soulé, C. See H. Gillet

Stancu, A. and E. Werner New higher-order equiaffine invariants  221

Struppa, D. C. See F. Colombo Sugawa, T. See S. Ruscheweyh Tenenblat, K. See R. Pina Tiep, P. H. See G. Navarro Veselý, L. See C. De Bernardi Vuorinen, M. See V. N. Dubinin Werner, E. See A. Stancu